OpenStudy (anonymous):

please help, ray BD is the bisector of angle ABC, find angle ABC angle CBD and angle DBA when angle ABD= 2x^2 and angle ABC=100

OpenStudy (anonymous):

Well, angle ABC is given and is the whole. Because of the bisection, angle ABD is half of the whole. angle CBD is congruent because it is also half.

OpenStudy (anonymous):

You don't even need to solve for x here, just find half of 100.

OpenStudy (campbell_st):

Ok.. so ABC = ABD + CBD then CBD = ABC - ABD CBD = 2x^2 since BD is a bisector so ABC = 2x^2 + 2x^2 or 100 = 4x^2 solve for x